OpenStudy (anonymous):

can someone please help me, i really have no idea how to do this. roses cost 15$ per dozen, carnations cost 8$ per dozen and sara bought a total of 17 dozen flowers and paid a total of 192$. how many roses did she buy?

OpenStudy (anonymous):

roses, r carnations, c r+c=17 15r+8c=192 c=17-r 15r+8(17-r)=192 solve for r

OpenStudy (anonymous):

thank you so much !!!!!

OpenStudy (anonymous):

fan me please

OpenStudy (anonymous):

how do i do that, i just signed up im new.,

OpenStudy (anonymous):

just click the become a fan right next to my name